### v1

# Explicit Visual Prompt Instruction Tuning

Abstract
Large Multimodal Models (LMMs) have made significant breakthroughs with the advancement of instruction tuning. However, while existing models can understand images and videos at a holistic level, they still struggle with instance-level understanding that requires a more nuanced comprehension and alignment. Instance-level understanding is crucial, as it focuses on the specific elements that we are most interested in. Excitingly, existing works find that the state-of-the-art LMMs exhibit strong instance understanding capabilities when provided with explicit visual cues. Motivated by this, we introduce an automated annotation pipeline assisted by GPT-4o to extract instance-level information from images and videos through explicit visual prompting for instance guidance. Building upon this pipeline, we proposed Inst-IT, a solution to enhance LMMs in Instance understanding via explicit visual prompt Instruction Tuning. Inst-IT consists of a benchmark to diagnose multimodal instance-level understanding, a large-scale instruction-tuning dataset, and a continuous instruction-tuning training paradigm to effectively enhance spatial-temporal instance understanding capabilities of existing LMMs. Experimental results show that, with the boost of Inst-IT, our models not only achieve outstanding performance on Inst-IT Bench but also demonstrate significant improvements across various generic image and video understanding benchmarks. This highlights that our dataset not only boosts instance-level understanding but also strengthens the overall capabilities of generic image and video comprehension.
Equal contributions. Corresponding author.
Figure 1: Current LMMs struggle with instance-level understanding, failing to capture the nuanced details about specific instances. To address this, we create a large-scale instance-specific instruction tuning dataset and train a multimodal model on it. Compared to existing models, our model show better performance in instance-level understanding.
1 Introduction
Recently, Large Multimodal Models (LMMs) have seen remarkable advancements. A key breakthrough is visual instruction tuning [39, 16], enabling models to follow any type of user instructions. This paves the way to building general-purpose multimodal assistants capable of handling a wide range of real-world tasks [30]. Inspired by this initial work, numerous follow-up studies have emerged in both image-language [40, 54, 99, 10, 12, 73] and video-language [46, 37, 90, 33, 72, 76, 74] domains. However, although these models can understand the entire images or videos, they still struggle to comprehend instance-specific content, as illustrated in Fig. 1. This hinders them from a fine-grained understanding of visual inputs.
Instance-level understanding refers to comprehending the attributes of specific instances within an image or video, as well as the relationships and interactions between them. This requires models to exhibit nuanced comprehension and fine-grained alignment. Instance understanding has been a long-standing pursuit of the computer vision community with extensive efforts devoted to object detection [65, 63, 52], instance segmentation [66, 53] and object tracking [18], etc. This capability is essential for real-world applications, where users pay more attention to the instances that they are interested in. In the era of LMMs, although there have been some attempts in exploring multimodal instance understanding [22, 5, 94, 98, 95], they are primarily limited in the image domain, leaving the videos under-explored. Compared to images, understanding instances in videos is considerably more challenging, as it requires not only capturing their spatial information but also their temporal dynamics. Driven by this, we aim to advance the multimodal instance-level understanding in both images and videos. To this end, we focus on three key aspects of the effort: the evaluation benchmark, the instruction-tuning dataset, and the training recipe.
We begin by introducing an automated data annotation pipeline designed to provide detailed annotations for images and videos, with a particular emphasis on individual instances. Specifically, this pipeline first highlights the instances of interest in the images and videos using instance-level visual prompts [81]. Then, GPT-4o [55] is leveraged to generate fine-grained and multi-level annotations that covering several aspects, including: instance-level captions, frame-level captions, temporal changes descriptions, video-level captions, and open-ended question-answer pairs.
With the help of this pipeline, we first carefully design the Inst-IT Bench, a benchmark for diagnosing multimodal models in instance-level understanding, which includes 1,000 image test samples and 1,000 video test samples. We conduct extensive evaluations on Inst-IT Bench, and find that existing open-source models show limited performance in understanding individual instances. To address this gap, we apply the propose data annotation pipeline to perform fine-grained annotations on nearly 21k videos and 51k images, resulting in Inst-IT Dataset, which contains a total of 207k frame-level annotations, 21k video-level annotations, and 335k open-ended question answer pairs. To the best of our knowledge, this is the first instruction-tuning dataset that provides fine-grained annotations focused on instance of interest in both images and videos. Based on Inst-IT Dataset, we propose a continuous instruction-tuning recipe that effectively combines our instance understanding datasets with other general instruction-tuning data. By incorporating this small amount of data, the enhanced models demonstrate strong performance not only on Inst-IT Bench, but also on several widely used image and video benchmarks, such as AI2D [25] and Egoschema [47].
Our contributions are three-fold:
  1. 1.
    An instance-centric understanding benchmark. We introduce Inst-IT Bench, a benchmark specifically designed to evaluate instance-level understanding in multimodal models, encompassing both image and video understanding, and perform extensive evaluations on it.
  2. 2.
    An instance-grounded instruction-tuning dataset. We develop Inst-IT Dataset, the first instruction-tuning dataset that includes both images and videos, featuring explicit instance-level visual prompts. It is built through a GPT-4o-assisted automated annotation pipeline.
  3. 3.
    An instance-enhanced Large Multimodal Model. We further integrate Inst-IT Dataset into the instruction tuning of LMMs. We propose a continuous instruction tuning approach to effectively enhance spatial-temporal instance understanding while consistently improving general comprehension.
2 Related Work
Large Multimodal Models. Recently, significant progress has been witnessed in large multimodal models (LMMs). BLIP-2 [32] and Flamingo [2] leverage visual resamplers to integrate image features as language inputs by extracting a fixed number of visual tokens. In a parallel line of research, LLaVA [39] and its follow-upds [54, 40, 29, 97, 38, 51, 91] have achieved remarkable success by connecting vision and language through a simple projection module. This approach has simplified the pipeline, even outperforming previous models with less training effort. Additionally, researchers are extending LMMs’ capabilities to temporal understanding by incorporating multi-frame inputs [37, 74] or explicit temporal modules [36, 23], bringing these models closer to applications in real-world scenarios. However, we observe that existing LMMs struggle with instance-level understanding and often fail to accurately follow instructions to ground specific instances. We emphasize the importance of instance understanding and enhance it through instruction fine-tuning with explicit visual prompts, which also enhances general understanding.
Multimodal Datasets and Benchmarks. With the rapid progress in LMMs, numerous instruction tuning datasets have been developed. LLaVA-Instruct [39] is one of the first to use GPT-4, leveraging object categories, bounding boxes, and image-level captions to generate diverse visual instruction tuning data. Building on this, follow-up studies have focused on enhancing large language and multi-modal models to generate synthetic data [10, 73, 8] or improve the annotation pipeline [11, 97]. Simultaneously, various multimodal benchmarks have been proposed to evaluate LMMs across different aspects, such as comprehensive understanding [28], OCR and text comprehension [48, 49, 50, 70], temporal understanding [21, 47, 77, 6, 42], hallucination [34], and instruction-following [60]. However, they focus more on image or video-level understanding and lack fine-grained emphasis on specific instances. In this work, we emphasize the importance of instance understanding in both images and videos. Specifically, we propose the Inst-IT Bench to evaluate the instance understanding of LMMs and create the Inst-IT Dataset, providing detailed instance-level annotations to enhance instance understanding capabilities.
Multimodal Instance Understanding. Understanding individual instances in visual inputs is a central focus in computer vision, with key tasks such as object detection [65, 63], instance segmentation [66, 27], and object tracking [18, 45, 84]. In the era of large multimodal models, instance-level understanding has also gained increasing attention. SPEC [56], ARO[88] and Winoground [69], evaluates the ability of vision-language models to comprehend fine-grained visual-linguistic concepts, revealing that models like CLIP [61] struggle to grasp attributes of specific instances or the relationships between them. To address this, works like KOSMOS-2 [57] and Shikra [9] encode instance information in textual form. In parallel, another line of research, including SoM-LLaVA [79], RegionGPT [22], GPT4ROI [94], MG-LLaVA [98], OMG-LLaVA [95], and ViP-LLaVA [5], explores the use of visual prompting to guide models in focusing on specific instances. Among these, SoM-LLaVA [79] is most closely related to our work. It uses SoM [81] to prompt images and asks models to list the instances present, finding that instance-level understanding enhances global understanding. However, SoM-LLaVA [79] focuses solely on the image domain and does not address the more challenging video domain. In contrast, we focus on both images and videos, aiming to advance multimodal models in understanding the spatiotemporal changes of individual instances.
Figure 2: The automated data generation pipeline. We process the video frames sequentially. At each timestamp $t$, GPT-4o is prompted to create a frame-level annotation $Y_{t}^{f}$ based on the current frame $X_{t}$ and the previous frame $X_{t\text{-}1}$. Then, all the frame-level annotations are aggregated to produce a video-level description $Y^{vid}$ and a set of open-ended question-answer pairs $Y^{qa}$.
Figure 3: Visualization of an data example from our Inst-IT Dataset. For each video, we provide (a) frame-level annotations, (b) a video-level description, and (c) open-ended question-answer pairs. Each frame-level annotation consists of three parts: captions for individual instances, a caption for the entire scene, and captions describing the temporal changes involving specific instances. A key feature of our dataset is its emphasis on instances of interest, including their state in each frame, how they change between frames, and questions and answers focused on their specific details throughout the video. The contours of instances in this example are deliberately highlighted for better visualization. A complete data example can be found in Sec. C.2.
3 Inst-IT
High-quality instruction following data are essential for training effective multimodal models. To address the shortage of instance-grounded data, we propose an automated pipeline to generate detailed annotations for both images and videos, with a special emphasis on Instances of Interest. We begin by collecting a diverse set of images and videos, construct visual prompts and leverage GPT-4o to generate instruction-following data based on these visual prompted inputs (Sec. 3.1). Building upon this pipeline, we carefully design a benchmark to assess models in instance-specific understanding (Sec. 3.2). Furthermore, we build a large-scale vision language dataset (Sec. 3.3) and proposed a continuous instruction-tuning pipeline (Sec. 3.4) to enhance the instance understanding capabilities of LMMs.
3.1 Instance-centric Annotation Pipeline
Overview. We propose an automated annotation pipeline designed to annotate videos with a particular focus on individual instances. As illustrated in Fig. 2, the overall process involves annotating each frame sequentially, combining these frame-level annotations into a comprehensive video-level description, and generating a set of open-ended questions about the video content.
Visual Prompting. To generate more fine-grained annotations focused on specific instances, we augment the images and videos with visual prompts to highlight the instances of interest. Specifically, we use the Set-of-Marks (SoM) visual prompting [81], which overlays a numerical ID directly on each instance. We find this method highly effective in guiding GPT-4o to provide annotations focused on individual instances. For more details on SoM prompting, please refer to Appendix A.
Frame-level Annotation. We annotate each frame of the video sequentially. At each timestamp $t$, we provide GPT-4o with the current frame $X_{t}$, the previous frame $X_{t\text{-}1}$, and a tailored task prompt $P^{f}$. GPT-4o then produces a frame-level annotation $Y_{t}^{f}\text{=}(y_{t}^{ins},y_{t}^{img},y_{t}^{dif})$ across three aspects, where $y_{t}^{ins}$ represents the captions for individual instances, $y_{t}^{img}$ is caption for the entire image, and $y_{t}^{dif}$ describes the temporal differences from the previous frame:
$Y_{t}^{f}=\texttt{GPT}(P^{f},X_{t},X_{t\text{-}1})$ (1)
Video-level Summary. After obtaining annotations for each frame, we employ GPT-4o to aggregate them into $Y^{vid}$, a comprehensive description for the entire video:
$Y^{vid}=\texttt{GPT}(P^{vid},[Y_{1},Y_{2},\cdots,Y_{N}]),$ (2)
where $P^{vid}$ is the task prompt designed for video-level summary and $N$ is the total number of frames.
Open-Ended Question Answer Pairs. Based on the frame-by-frame annotations, we also prompt GPT-4o with the task prompt $P^{qa}$ to create $M$ open-ended QA pairs $Y^{qa}\text{=}\{(q_{i},a_{i})\}_{i=1}^{M}$ focusing on specific instances:
$Y^{qa}=\texttt{GPT}(P^{qa},[Y_{1},Y_{2},\cdots,Y_{N}])$ (3)
Putting It Altogether. Through the steps outlined above, we can annotate the video content at multiple levels of granularity, highlighting the elements related to specific instances. Specifically, as illustrated in Fig. 3, each video is accompanied by the following three types of annotations:
  • $N$ frame-level annotations, each encompassing a detailed description of individual instances, the entire image, and temporal changes.
  • A comprehensive description covering the entire video.
  • $M$ open-ended question-answer pairs that focused on specific instances or the inter-instance relationships.
Additional information about the design of each task prompts, as well as the organization of the input for GPT-4o is provided in Appendix A.
3.2 Inst-IT Bench
Existing multimodal benchmarks primarily focus on global understanding, failing to provide more in-depth insights into the instance-level comprehension capability of models. To fill this gap, we present Inst-IT Bench, a benchmark specifically designed to diagnose instance-level understanding in both images and videos.
Construction Process. Inst-IT Bench is built through a semi-automated workflow. Initially, we apply the pipeline described in Sec. 3.1 to generate around 20 open-ended QA pairs for each image and video. These generated pairs are then manually reviewed to ensure both accuracy and diversity. Specifically, we remove overly simple questions, ensuring the remaining ones are focused on specific instances. We also refine the questions and answers, making necessary rephrasing to ensure correctness. After this rigorous process, each sample retains 2 to 5 carefully selected question-answer pairs from the original 20, forming the final test set. In addition, we generate three hard negative answers for each question, forming a multiple-choice QA with four options. In total, Inst-IT Bench includes 1,000 QA pairs for 338 images and 1,000 QA pairs for 206 videos. Each QA pair is available in both open-ended and multiple-choice formats. More details can be found in Appendix B.
Metrics. Each test sample in Inst-IT Bench includes both open-ended and multiple-choice formats. For open-ended QAs, we leverage GPT-4o to score the response from a model based on its similarity to the ground-truth answer. For multiple-choice QAs, we calculate the average accuracy across all samples. More details on the metric calculations can be found in Appendix B.
3.3 Inst-IT Dataset
Instruction fine-tuning is crucial for training multimodal models, however, we find there lacks instruction fine-tuning datasets that focus on specific instances. This gap hinders the development of instance-level understanding of the community. Using the data annotation pipeline presented in Sec. 3.1, we create a large-scale instruction-tuning dataset Inst-IT Dataset. To the best of our knowledge, this is the first instruction fine-tuning dataset that provides fine-grained annotations centric on specific instances
Data Sources. We utilize five video instance segmentation datasets (BRUST [3, 17], UVO [75], OVIS [59], LVVIS [71] and YoutubeVIS-2021 [82]) and two object tracking datasets (BenSMOT [35], VidOR [67]) as our video sources, as they provide annotations of instance locations, which is useful in SoM visual prompting [81]. For the image source, we select the SA-1B [27] dataset due to its diversity and abundance of instance objects. Altogether, we collect 51k images and 21k video samples as our visual data pool. More details can be found in Appendix C.
Statistics. Inst-IT Dataset contains 21k videos and 51k images. On average, each video includes one video-level description, 7.3 frame-level descriptions (each covering individual instances, the entire scene, and temporal differences), and 15.6 open-ended question-answer pairs. We treat images as static, single-frame videos without temporal descriptions. In total, Inst-IT Dataset includes visual data from 21k videos and 51k images, along with 21k video-level descriptions, 207k frame-level descriptions, and 335k open-ended QA pairs. For more details about Inst-IT Dataset, please refer to Appendix C.
3.4 Instruction Tuning with Inst-IT Dataset
We adopt the widely-used LLaVA-Next [41] architecture to evaluate the effectiveness of our Inst-IT Dataset. We train our model under an image-video joint training pipeline, where we mix our Inst-IT Dataset with the open-source LLaVA-Next-DATA [43]. For single-image samples, we follow the original AnyRes paradigm [41] to split and encode sub-images according to the aspect ratio. For video and multi-image data, we batch the samples together, encode them, and flatten them into a sequence. Additionally, we apply $2\times 2$ spatial pooling to reduce the number of visual tokens in the video inputs. More details are in Sec. 4.1.
Model LLM #IT Image Video Open-Ended Q&A Multi-Choice Q&A Open-Ended Q&A Multi-Choice Q&A Random Guess - N/A - 25.0 - 25.0 GPT-4o [55] - N/A 74.1 84.8 65.5 81.0 Gemini-1.5-flash [64] - N/A 65.3 79.5 57.9 75.8 Open-source image models LLaVA-1.5 [40] Vicuna-7B [15] 665K 41.6 32.1 - - ViP-LLaVA [5] Vicuna-7B [15] $\sim$1.2M 42.1 29.2 - - SoM-LLaVA [79] Vicuna-7B [15] 695K 45.1 40.0 - - LLaVA-Next [41] Vicuna-7B [15] 765K 46.0 42.4 - - Open-source video models LLaVA-NeXT-Video [96] Vicuna-7B [15] 860K 46.5 39.5 25.8 24.8 ShareGPT4Video [11] Llama3-8B [19] $\sim$1.0M 43.2 48.7 27.8 16.1 MiniCPM-V 2.6 [83] Qwen2-7B [80] $\sim$7.0M 57.6 66.8 40.0 45.2 LLaVA-OV (SI) [29] Qwen2-7B [80] $\sim$7.2M 60.3 61.8 31.4 36.4 LLaVA-OV [29] Qwen2-7B [80] $\sim$8.8M 48.0 71.7 33.2 45.6 LLaVA-Video [97] Qwen2-7B [80] $\sim$7.4M 45.1 67.0 34.1 53.2 InternVL2 [12] InternLM2.5-7B [7] N/A 58.6 66.5 39.8 45.5 Qwen2-VL-Instruct [74] Qwen2-7B [80] N/A 48.3 64.9 38.2 59.4 Qwen2-VL-Instruct [74] Qwen2-72B [80] N/A 55.5 74.7 45.5 74.6 Our models LLaVA-Next-Inst-IT Vicuna-7B [15] 920K 68.6 63.0 49.3 42.1 LLaVA-Next-Inst-IT Qwen2-7B [80] 920K 67.9 75.3 45.7 53.3
Table 1: Performance of LMMs on Inst-IT Bench. We conduct extensive evaluations on Inst-IT Bench, including state-of-the-art open-source image models, video models, and cutting-edge proprietary models. #IT indicates the number of training samples used during the instruction-tuning stage. N/A indicates that the number of training samples is unknown.
4 Experiments
4.1 Implementation Details
We use LLaVA-Next [41] as our baseline multi-modal large model due to its widespread adoption. In the default configuration, Vicuna-1.5-7B [15] serves as the language model with CLIP-ViT-336 [62] as the vision encoder. We utilize the AdamW [44] with a cosine learning rate schedule for optimization. During the vision-language alignment stage, we use the LCS-558K dataset [40], and for the supervised fine-tuning stage, we leverage the open-source LLaVA-Next-DATA [43]. For single images, we split the original image into up to 4 sub-images based on its aspect ratio following the AnyRes [41] approach, and then concatenate the global image with these sub-images. For multiple images and video inputs, we skip the AnyRes procedure and encode every single image. Additionally, we apply $2\times 2$ spatial pooling to reduce the number of visual tokens for video inputs. We limit the maximum number of frames to 32 and the context length of LLMs to 6K due to GPU and memory constraints. To enhance instance-level understanding with our Inst-IT Dataset, we combine Inst-IT Dataset with LLaVA-Next-DATA and introduce an additional continuous supervised fine-tuning (SFT) stage. In this stage, we freeze the first 12 layers of the vision encoder to mitigate potential distribution shifts caused by visually prompted images. Furthermore, we use Qwen2-7B [80] with SigLIP-SO400M-384 [89] for improved performance in our main experiment, and Qwen2-1.5B with CLIP-ViT-336 for efficiency in our ablation study.
Method LLM Vision Encoder AI2D [25] MMMU [87] POPE [34] GQA [24] MM-Vet [85]
(test) (val) (test F1) (val) (test)
LLaVA-1.5 [40] Vicuna-7B [15] CLIP-ViT-Large [61] 54.8 35.3 85.9 62.0 30.5
LLaVA-Next [41] Vicuna-7B [15] CLIP-ViT-Large [61] 66.6 35.1 86.4 64.2 44.1
DeepStack-L [54] Vicuna-7B [15] CLIP-ViT-Large [61] - 35.7 86.7 63.1 29.9
DeepStack-L-HD [54] Vicuna-7B [15] CLIP-ViT-Large [61] - 35.6 86.5 65.2 37.5
VILA [38] Vicuna-7B [15] CLIP-ViT-Large [61] - - 85.5 62.3 34.9
ShareGPT4V [10] Vicuna-7B [15] CLIP-ViT-Large [61] - - - 37.6
MM-1.5 [91] MM-LLM-7B [91] MM-CLIP [91] 72.0 41.8 88.6 - 42.2
InternVL2 [12] InternLM-7B [68] InternViT-300M [12] 83.8 49.3 - - 60.0
LLaVA-OV (SI) [29] Qwen2-7B [80] SigLIP-SO400M [89] 81.6 47.3 - - 58.8
LLaVA-OV [29] Qwen2-7B [80] SigLIP-SO400M [89] 81.4 48.8 - - 57.5
Qwen2-VL-Instruct [74] Qwen2-7B [80] DFN-CLIP-H [20] 83.0 54.1 - - 62.0
LLaVA-Next-Inst-IT Vicuna-7B [15] CLIP-ViT-Large [61] 71.0 37.4 87.2 65.9 38.1
LLaVA-Next-Inst-IT Qwen2-7B [80] SigLIP-SO400 [89] 78.7 42.7 87.6 65.5 44.7
Table 2: Main results on image benchmarks.
Method LLM Vision Encoder ANet-QA [86] EgoSchema [47] Next-QA [77] VideoMME [21] TempCompass [42]
(open-ended) (subset) (val) (w/o subs) (3 avg)
DeepStack-L [54] Vicuna-7B [15] CLIP-ViT-Large [61] 49.3 38.4 61.0 - -
IG-VLM [26] Vicuna-7B [15] CLIP-ViT-Large [61] 54.3 35.8 63.1 - -
LLaVA-Next [41] Vicuna-7B [15] CLIP-ViT-Large [61] 53.8 - - - -
SF-LLaVA [78] Vicuna-7B [15] CLIP-ViT-Large [61] 55.5 47.2 64.2 -
Video-ChatGPT [46] Vicuna-7B [15] CLIP-ViT-Large [61] 35.2 47.3 - - -
VideoLLaMA2 [13] Vicuna-7B [15] CLIP-ViT-Large [61] 50.2 - 51.7 - -
LLaVA-Next-Video [96] Vicuna-7B [15] CLIP-ViT-Large [61] 53.5 43.9 - 46.5 -
LLaVA-Next-Video-DPO [96] Vicuna-7B [15] CLIP-ViT-Large [61] 60.2 - - 58.3
LongVA [93] Qwen2-7B [80] CLIP-ViT-Large [61] - 68.3 52.4 61.3
MM-1.5-Video-SFT [91] MM-LLM-7B [91] MM-CLIP [91] 60.9 57.2 76.8 53.5 -
InternVL2 [12] InternLM-7B [68] InternViT-300M [12] - - - 54.0 -
LLaVA-OV [29] Qwen2-7B [80] SigLIP-SO400M [89] 56.6 60.1 79.4 58.2 69.4
LLaVA-Video [97] Qwen2-7B [80] SigLIP-SO400M [89] 56.5 57.3 83.2 63.3 -
Qwen2-VL-Instruct [74] Qwen2-7B [80] DFN-CLIP-H [20] - 66.7 - 63.3 72.9
LLaVA-Next-Inst-IT Vicuna-7B [15] CLIP-ViT-Large [61] 53.7 57.8 70.2 44.3 59.8
LLaVA-Next-Inst-IT Qwen2-7B [80] SigLIP-SO400 [89] 55.2 50.4 73.0 54.0 63.9
Table 3: Main results on video benchmarks. We report the average of 3 parts (MCQA, Y/N, Caption Match) of TempCompass for determinism results.
4.2 Main Experiments
Results on Inst-IT Bench. We conduct extensive evaluate of existing multimodal models on Inst-IT Bench. The results in Tab.1, show that with instruction tuning on LLaVA-NeXT [41] using Inst-IT Dataset, the model achieves a significant improvement of nearly 20% on average score, validating the effectiveness of our Inst-IT Dataset. Moreover, although ViP-LLaVA [5] utilizes visual prompts for instruction tuning, it shows minor improvement over the baseline LLaVA-1.5 [40], possibly due to overfitting to its training data. In contrast, our model demonstrates consistent improvements on other instance understanding benchmarks, such as ViP-Bench [5], as well as on general-purpose evaluation sets like AI2D and Egoschema (further discussions are in the following sections). This suggests that model trained with Inst-IT Dataset can generalizes well to out-of-domain data. Besides, as Qwen2VL-72B does not show substantial improvements over its smaller 7B model, indicating that simply scaling up the model size cannot address the challenges in instance understanding. Similarly, by comparing the amount of instruction tuning data used by each models, we observe that large-scale coarse-grained annotations does not lead to essential improvements either. This highlights the importance of instance-specific annotated data, as we achieve impressive performance by training a 7B model on 920k data.
Results on academic benchmarks. We evaluate our models using the LMMs-Eval [92] toolkit. To ensure a fair comparison with other models, we primarily report results from their original papers or from reproduced results in other studies. On widely-used image benchmarks shown in Tab. 2, our Inst-IT consistently outperforms LLaVA-Next, our direct baseline model. Furthermore, when utilizing a more advanced language model and a vision encoder, our method achieves performance comparable to large-scale SFT LMMs, such as LLaVA-OV and Qwen2-VL-Instruct, while requiring significantly less computational and data cost for instruction tuning. For video understanding benchmarks Tab. 3, our method significantly outperforms both LLaVA-Next and LLaVA-Next-Video. These consistent improvements demonstrate that enhancing instance-level understanding through explicit visual prompt instruction tuning is an effective strategy for improving generic spatiotemporal understanding.
We further evaluate our model on ViP-Bench [5], an instance-level understanding benchmark that aligns closely with the objectives of our Inst-IT. However, while ViP-Bench focuses primarily on domain-specific instance-level referring comprehension, our Inst-IT encompasses a broader scope, involving cross-instance and cross-timestamp interactions, among other aspects. As shown in Tab. 6, our model demonstrates strong generalization capabilities when evaluated on downstream tasks. Specifically, our Inst-IT with Vicuna-7B achieves comparable performance to ViP-LLaVA when using rectangular bounding boxes as visual prompts and even surpasses ViP-LLaVA when employing human-style visual prompts. Notably, our model operates as a generalist under zero-shot evaluation, whereas ViP-LLaVA benefits from in-domain tuning since it is fine-tuned on the in-domain referring comprehension dataset.
CL Tune Enc Data comb AI2D MMMU POPE GQA Inst-IT-I Next-QA VideoMME Inst-IT-V
(test) (val) (test F1) (val) (mc) (mc) (w/o subt) (mc)
All layers LLaVA-Next 61.1 35.9 86.9 61.4 45.3 56.6 45.7 31.3
All layers LLaVA-Next & Inst-IT Dataset video 60.7 34.7 86.1 61.2 60.7 59.7 47.1 43.0
All layers LLaVA-Next & Inst-IT Dataset video 62.3 35.7 86.7 62.9 61.8 62.4 46.7 44.4
None LLaVA-Next & Inst-IT Dataset video 63.1 35.0 86.9 62.5 60.2 63.2 47.2 44.3
Last 12 LLaVA-Next & Inst-IT Dataset video 63.2 34.9 87.0 62.5 60.1 63.3 47.2 44.0
Last 12 LLaVA-Next & Inst-IT Dataset (img+vid) 63.0 36.1 87.2 62.7 59.6 64.3 46.6 43.7
Table 4: Ablation on data training recipe. We utilize LLaVA-Qwen2-1.5B-CLIP336 as the baseline model. Inst-IT-I and Inst-IT-V indicate the multi-choice splits of the image and video part of our Inst-IT Bench, separately.
# Data Combination AI2D MMMU POPE GQA Inst-IT-I Next-QA VideoMME Inst-IT-V
(test) (val) (F1) (val) (mc) (mc) (w/o subt) (mc)
0 LLaVA-Next 61.1 35.9 86.9 61.4 45.3 56.6 45.7 31.3
1     $+$ inst-cap & img-cap 63.0 35.1 86.1 62.7 58.9 62.4 46.0 33.8
2     $+$ temporal diff 63.0 35.6 87.1 62.7 59.6 64.2 45.6 36.9
3     $+$ video-description & qa 63.2 34.9 87.0 62.5 60.1 63.3 47.2 44.0
4     $+$ Inst-IT Dataset image 63.0 36.1 87.2 62.7 59.6 64.3 46.6 43.7
Table 5: Ablation on detailed data combination. The dataset combination in line #3 corresponds to the video part of Inst-IT Dataset, while line #4 represents the complete Inst-IT Dataset by incorporating the image part into line #3.
Model Synthesized visual prompts (tight bounding box) Visual prompts from human (e.g. arrow, circle)
Rec OCR Know Math Rel Lang All Rec OCR Know Math Rel Lang All
GPT-4V-turbo-detail:high [1] 58.1 69.8 59.5 71.0 61.4 51.9 60.7 56.9 69.7 63.7 80.6 61.1 45.6 59.9
GPT-4V-turbo-detail:low [1] 53.2 50.3 55.6 67.7 57.5 57.5 52.8 51.7 50.3 59.3 60.3 55.0 43.8 51.4
InstructBLIP-7B [16] 36.9 16.3 34.2 22.3 26.8 7.5 31.7 38.9 17 35.4 9.7 29.3 17.5 33.3
Shikra-7B  [9] 40.2 10.0 28.0 3.5 18.9 20.6 33.7
GPT4ROI-7B  [94] 35.6 16.7 29.7 9.7 32.5 13.8 35.1
Kosmos-2 [58] 29.5 14.2 18.5 9.7 7.5 21.9 26.9
LLaVA-1.5-7B [40] 50.8 12.4 49.2 6.5 51.8 23.8 41.6 49.1 13.0 42.9 9.7 50.0 27.5 40.2
Qwen-VL-Chat [4] 43.0 30.4 40.2 9.7 25.7 28.7 39.2 48.7 22.1 41.2 6.5 48.2 25.0 41.7
ViP-LLaVA-7B [5] 54.8 18.8 52.9 9.7 53.9 42.5 45.5 55.3 17.6 45.9 8.1 44.6 33.1 46.8
LLaVA-Next-Inst-IT-Vicuna-7B 51.3 23.7 54.2 12.9 64.3 46.2 45.1 55.0 21.3 52.5 16.1 57.5 40.6 48.2
LLaVA-Next-Inst-IT-Qwen2-7B 58.9 24.5 48.5 12.9 48.2 46.3 50.5 57.7 22.5 53.2 19.4 53.6 45.0 49.0
Table 6: ViP-Bench Evaluation Results. We direct perform evaluation with our Inst-IT models.
4.3 Ablation Study
We use Qwen2-1.5B [80] as the language model and CLIP-ViT-L-336 [61] as the vision encoder for all our ablation experiments, prioritizing efficiency. We first conduct an ablation study on the training recipe to investigate how to effectively integrate Inst-IT Dataset with existing academic SFT datasets [43] for a balanced improvement. Next, we perform a detailed analysis of the impact of each component in our Inst-IT Dataset.
Effectiveness of our continuous instruction-tuning paradigm. As shown in Tab. 4, directly mixing the video Inst-IT Dataset with LLaVA-Next-DATA leads to significant improvements on video benchmarks. However, the performance on generic image understanding slightly declines. We believe this is due to two main reasons: (1) the increased ratio of video data may suppress image understanding; (2) visually prompted images may introduce a distribution shift from natural images. To address these issues, we propose a continuous SFT paradigm based on single-image models and freeze the first 12 layers of the vision encoder to preserve realistic low-level features. As a result, our model achieves balanced performance with this training approach.
Detailed dataset combination. To further investigate the effectiveness of each component in Inst-IT Dataset, as illustrated in Fig. 3, we conduct an extensive ablation study by progressively adding data components. As shown in Tab. 5, the instance-level and image-level frame captions are essential for improving instance understanding in images. Meanwhile, temporal differences, along with video-level descriptions and QA, significantly enhance video instance understanding. Finally, incorporating the image component of Inst-IT enables our model to achieve the best-balanced performance across generic image and video understanding benchmarks, as well as our proposed instance understanding benchmark.
5 Conclusion
Instance-level understanding that detects, segments, and reasons nuanced relationships among objects has long been the goal of computer vision research, yet limited effort has been made to equip Large Multimodal Models with such capabilities. In this paper, we introduced Inst-IT Bench, a carefully curated benchmark for evaluating multimodal instance understanding abilities. Extensive evaluations for a wide range of open-source models demonstrate the limitedness of current models for understanding at the instance level. To mitigate this issue, we collected Inst-IT Dataset, the first instruction-tuning dataset with explicit instance-level visual prompting and annotation. Building upon Inst-IT Dataset, we proposed Inst-IT, a continuously fined-tuned framework that excels in instance understanding as well as widely used image and video benchmarks.
Limitation and Future Works. Our current experiments are conducted on 7B and 1.5B models due to the computation cost. Moreover, our current data pipeline is automated but constrained by the overhead of GPT-4o. We can further scale the model size and scale the dataset using a model-in-the-loop approach and improve the model through multi-round instruction tuning with self-synthesized data. We leave this direction for future work.
Appendix A Data Annotation Pipeline
A.1 Set-of-Marks Visual Prompting
Performing instance-level annotations is challenging, and we adopt the SoM visual prompting technique [81] to address this. Specifically, as illustrated in Fig. 4, we overlay a numeric ID at the center of each instance and maintain the same ID for a given instance across all frames. This simple augmentation can explicitly guide GPT-4o to focus more effectively on the instances of interest, enabling finer-grained and more accurate annotations. Furthermore, segmentation masks are necessary to calculate the center coordinates of each instance. Details on how these masks are obtained are provided in Sec. C.1.
A.2 Prompting GPT-4o
Task prompt templates. Prompt engineering is crucial for enabling GPT-4o to accomplish specific tasks. In this section, we present the task prompts that we designed to prompt GPT-4o for data annotation:
  • The task prompt $P^{f}$ for frame-level annotation, Fig. 6.
  • The task prompt $P^{vid}$ for video-level annotation, Fig. 7.
  • The task prompt $P^{qa}$ for open-ended question-answer pairs generating, Fig. 8.
GPT-4o API version. During the annotation process, we use the GPT-4o-2024-08-06 API and leverage its structured output functionality to facilitate output parsing, enabling the model to respond in a predefined JSON format.
Figure 4: Set-of-Marks visual prompting on the original videos. Each instance is assigned a unique numeric ID, which remains consistent across all frames.
Figure 5: GPT-4o-based open-ended question answering correctness assessment. The underlined parts in the figure are included only when evaluating the video split, while the italicized parts will be replaced by the actual sample for scoring.
Figure 6: Frame-level annotation task prompt, the italicized part are placeholders for the actual inputs.
Figure 7: Video-level annotation task prompt, the italicized part are placeholders for the actual inputs.
Figure 8: Open-ended question-answer pairs generation task prompt, the italicized part are placeholders for the actual inputs.
Figure 9: A data example from Inst-IT Bench. Each test sample includes both open-ended QA and multiple-choice QA, focusing on specific instances or the relationships and interactions between instances.
Appendix B More Details about Inst-IT Bench
B.1 Negative Options Generation
We use the ground-truth from open-ended QA as the positive option and additionally craft three negative options, forming a multiple-choice question with four options. To create hard negatives, we first have the model answer the open-ended questions and use GPT-4o to score the correctness of the responses. If the score is lower than 0.4, we consider it a difficult negative answer and include it as one of the negative options. Finally, we randomly shuffle the four options to ensure that the correct one appears in each position with equal probability.
B.2 LLM-based Evaluator for Open-Ended QA
Recent studies [85, 14] suggest that LLMs can serve as effective evaluators. Building on this, we use GPT-4o to assess the accuracy of open-ended question answering. Specifically, GPT-4o assigns a score between 0 and 1 based on three key factors: the question, the ground-truth answer, and the model prediction. Given that Inst-IT Bench prioritizes instance-level understanding, we pay special attention to the accuracy of instance ID references. Furthermore, for the video split of Inst-IT Bench, we emphasize the correctness of timestamps to ensure temporal correctness. The task prompt for GPT-4o is illustrated in Fig. 5.
B.3 Data Example
To provide a clearer understanding of Inst-IT Bench, we present a data example in Fig. 9. Each question includes both open-ended and multiple-choice formats, focusing on specific instances or exploring the relationships and interactions between multiple instances. This design highlights the significant distinction from other benchmarks, emphasizing fine-grained understanding at the instance level.
Appendix C More Details about Inst-IT Dataset
C.1 Data Collection and Processing
Collection. We select five instance segmentation datasets and two multi-object tracking datasets as sources of video data. To prevent data leakage, we only used the training splits of these datasets. Additionally, we use the SA-1B [27] dataset as a source of image data, and only utilize the first five data splits provided by officially. In total, we collect 21,418 videos and 51,101 images. Tab. 7 provides detailed statistics on the sources of our data.
Processing. When constructing SoM [81] visual prompts, we need to obtain the mask annotations for each instance to determine the location of the numeric IDs. For the video instance segmentation datasets [3, 75, 59, 71, 82], the instance masks are already provided and can be used directly. For multi-object tracking datasets [35, 67], we use their bounding box annotations as prompts and generate instance masks through the SAM [27]. For images in the SA-1B dataset [27], we employ Semantic-SAM [31] to segment the instances and obtain their masks.
C.2 Data Example
In this section, we provide a complete video annotation from Inst-IT Dataset to offer a clearer understanding of its content and format. In all annotations, we use the format [ID] to refer to instances and $<$timestamp$>$ to refer to timestamps. We present the frame-level annotations in Tab. 8 to Tab. 10. We can see that each frame-level annotation $Y^{f}$ consists of three parts: instance-level descriptions $y^{ins}$, image-level descriptions $y^{img}$, and temporal differences $y_{dif}$. Additionally, each video is accompanied by a series of open-ended question-answer pairs $Y^{qa}$, most of which center on specific instances or their relationships, as illustrated in Tab. 11. Furthermore, we generate a dense video-level caption $Y^{vid}$ summarizing the entire video, as shown below:
Dataset Name Ann. Type Split Sample Num.
Video Instance Segmentation
BRUST [3] mask training 500
UVO [75] mask training 5,135
OVIS [59] mask training 599
LVVIS [71] mask training 3,057
YoutubeVIS [82] mask training 2,897
Video Object Tracking
BenSMOT [35] box training 2,261
VidOR [67] box training 6,969
Image
SA-1B [59] none 1-5 51,101
Table 7: Data sources. We use 7 video datasets and 1 image dataset as our data sources. We demonstrate their annotation formats, splits we used, and the number of samples from each dataset.
Video-level caption: The video appears to document a formal or celebratory event indoors, possibly a ceremony such as a wedding or official gathering, occurring in a room with wooden paneling and art or framed pictures on the wall. At the beginning, during $<$1$>$, [1] is wearing a light gray suit and stands with [2] in a sleeveless white lace dress, and [3] in a dark floral-patterned dress. The three are close together, suggesting an intimate or focused setting. The progression between $<$2$>$ and $<$3$>$ involves subtle changes in posture and interaction. [1] moves closer to [2], appearing to hold hands or engage in an exchange, possibly involving a ring, as indicated by a bouquet of flowers. [3] remains supportive and smiling, while [5], in red, momentarily holds an object above their head, before disappearing from view by $<$4$>$.In frames $<$5$>$ to $<$7$>$, [1] and [2] maintain a close interaction, suggestive of a significant moment such as an exchange of vows or rings. They are closely observed by [3], who stands smiling nearby, while [1] and [2] occasionally adjust their positions, facing each other initially and then turning outward, which may signal transitioning from an intimate moment to posing for a photo. By $<$7$>$, [4] joins, dressed in darker attire, emphasizing the formal setting as [3] is no longer visible.Through $<$8$>$ and $<$9$>$, the group dynamics change slightly with the absence of [4] and [3] entering the scene again. [1] and [2] appear to engage in a warm interaction as [3] supports them, clapping, alongside visible hands of [4] indicating applause, marking a cheerful tone.Finally, during $<$10$>$ to $<$12$>$, the focus shifts as [1] and [2] first engage in a kiss, underscoring an intimate conclusion to their ceremony. They later stand apart slightly at the center, with [1] smiling or speaking, and [2] leaning towards [1] suggestively content. Throughout, the consistent joyous mood is accentuated by [3]’s ongoing clapping and expression of joy, emphasizing shared celebration and approval from the audience captured.”
Frame Instance-level Captions Image-level Captions Temporal Differences
[Uncaptioned image] timestamp$<$1$>$ 1: Wearing a light gray suit with a white shirt, standing indoors. 2: Wearing a sleeveless white lace dress, holding an object in the hand. 3: Wearing a dark floral-patterned dress with long wavy hair. [1] [2] [3] are standing closely together in an indoor setting. [1] is on the left side wearing a formal, light gray suit with a white shirt. [2], in the middle, is wearing a sleeveless white lace dress, holding something in their hand. [3] is on the right side in a dark floral-patterned dress with long, wavy hair. They appear to be in a room with wooden paneling and some framed art on the wall. null
[Uncaptioned image] timestamp$<$2$>$ 1: A person wearing a gray suit with a white shirt, short hair. 2: A person in a white, sleeveless dress with long dark hair. 3: A person wearing a dark floral dress with long dark hair. 5: A person wearing red, partially visible in the background. 6: A small black cellphone held in a hand. The scene appears to be in an office setting with a wooden table at the foreground. [1] is standing to the left, facing [2], and appears to be holding [2]’s finger or hand. [2] stands slightly to the right, returning focus with [1]. [3] is to the right of [2], slightly in the background, smiling and looking forward. A bouquet of white flowers lies on the table near [2]. [5] is partially visible in the background on the right, seated and wearing red. [6] is a cellphone held by [5]. Background shows a wooden wall and a reflection in a window. [1] has moved closer to [2] and is now in contact with [2]’s hand. [2] has turned slightly towards [1] compared to the previous frame. [3] remains in a similar position, but the expression suggests more engagement with the scene. [5] and [6] have appeared in the frame; [5] is visible in the background holding [6]. The table with a bouquet of flowers is now visible, indicating a shift in camera angle slightly to include more of the right side of the room.
[Uncaptioned image] timestamp$<$3$>$ 1: Wearing a grey suit, standing beside [2] and slightly turned towards them. 2: Wearing a white, sleeveless dress with floral textures. Holding a bouquet of white flowers. 3: Wearing a dark patterned dress, standing slightly behind [2]. 4: Partially visible, wearing dark clothing, located at the edge of the left side of the frame. 5: Seated, wearing a red outfit. Holding a white object above their head, possibly obscuring their face. The scene shows [1] [2] [3] near a wooden conference table in a professional setting, possibly an office. [1] wears a grey suit and is standing to the left, engaged with [2] who is wearing a white dress and holding flowers. [3], who is in a patterned dress, stands closely behind [2]. The newly appeared [4] is seated to the far left, partially visible at the edge of the frame. [5] is seated on the right side, holding an object above their head, possibly obscuring their face. The room has wooden walls and a framed picture hanging on the wall. Object [5] has lifted an object above their head, possibly a piece of paper. Object [4] has appeared in the scene, seated on the left side of the frame, which was not visible earlier. The positions of objects [1], [2], and [3] remain unchanged, as does the background and setting of the room. Overall, no significant movement is noticed in terms of camera angle or position for objects [1] [2] [3].
[Uncaptioned image] timestamp$<$4$>$ 1: Wearing a light gray suit jacket, white dress shirt, and dark pants. 2: Wearing a white dress with a lace overlay, fitted at the waist. 3: Wearing a patterned dress with a floral design, strapless. 4: Visible part of a person wearing a dark shirt, seated or standing near the table. The setting appears to be indoors, with [1] [2] and [3] standing together around a table with a bouquet of flowers on it. [1] is interacting with [2], who is at the center, and they are possibly holding hands or engaged in some form of exchange. [3] is standing beside [2] and looking on, slightly leaning towards her. The room has wooden walls and a large framed picture in the background. The setting suggests a formal or ceremonial atmosphere, possibly a wedding or an official gathering. The camera angle is focused on this group, highlighting their interaction. [1] has moved slightly closer to [2], and they appear to be holding hands or exchanging something. [5] is no longer visible in the frame, possibly due to a change in camera angle or positioning of the individuals.
[Uncaptioned image] timestamp$<$5$>$ 1: An adult wearing a light gray suit with button details and a white shirt. The expression and stance suggest focus and engagement. 2: An adult in a white, lacy dress with thin straps. The person has long dark hair and appears to be smiling, holding hands with [1]. 3: An adult wearing a multicolored, patterned dress. The person has long, wavy hair and is smiling while observing [1] and [2]. The current frame captures a moment in an interior setting with [1] wearing a light gray suit, [2] in a white lace dress, and [3] in a patterned dress. [1] and [2] are engaged, with [1] facing [2] and holding their hand, suggesting an exchange, possibly a ring. [2] smiles, indicating a moment of happiness. [3] stands to the right, smiling and observing the interaction, detached but engaged with the scene. The background shows a wooden wall and framed picture, reflecting a formal environment possibly used for ceremonies. A bouquet of flowers rests on the table in front of the group. Between the previous and the current frame, [1] and [2] have shifted slightly closer, with [1] now directly holding [2]’s hand, indicating a progression in their interaction, possibly the continuation or conclusion of an exchange, such as the placing of a ring. [3] remains in a similar position but continues to observe [1] and [2], emphasizing their passive role in the interaction. There is no notable change in the background or environment.
Table 8: Inst-IT Dataset Frame-level Annotation, Part I (frame 1-5). Please zoom in to view the instance ID labels.
Frame Instance-level Captions Image-level Captions Temporal Differences
[Uncaptioned image] timestamp$<$6$>$ 1: [1] is wearing a grey suit with a white shirt, looking forward, standing upright and smiling slightly. 2: [2] is wearing a white sleeveless dress, with hair tied back, and is standing with a calm expression. 3: [3] is wearing a floral dress with an energetic expression, standing with arms slightly bent. The image depicts a formal setting with a group of three adults, [1], [2], and [3], standing closely together. The background features a wooden paneled wall and a framed picture. [1] and [2] are positioned in the center, both facing forward, suggesting they are the focus of the occasion. [1] is on the left, wearing a grey suit, and [2] is to the right of [1] in a white dress. They appear to be engaged in a ceremony or formal event. [3] is to the right of [2], wearing a floral dress, and displays a cheerful demeanor. The lighting is bright, illuminating their faces and creating a formal, celebratory atmosphere. Between the frames, there is a noticeable shift in the poses and expressions of [1] and [2]. In the current frame, [1] is now standing upright with a slight smile, while previously [1] was leaning towards [2], holding [2]’s hand, suggesting a shift from interaction to posing. [2], who was previously looking at [1], is now facing forward with a calm expression, indicating a change from an interactive pose to a more neutral one. Both [1] and [2] have adjusted their posture to face the camera more directly. [3] remains in similar positioning as before but has moved slightly closer to [2] and is displaying a more energetic expression, emphasizing the cheerful atmosphere. The objects on the table in the foreground, visible in the previous frame, are no longer the focal point, showing that the primary focus is now the individuals standing together.
[Uncaptioned image] timestamp$<$7$>$ 1: [1] is dressed in a grey suit with a white shirt, looking formal and neat. 2: [2] is wearing a white, sleeveless dress with a lightly patterned texture. 4: [4] is dressed in a dark outfit, including a dark scarf or similar accessory. In the current frame, [1] is positioned in the center, wearing a grey suit and a white shirt. [2] is to the right of [1], dressed in a white sleeveless dress. [4] appears on the left side of the image, wearing a dark outfit, which includes a scarf, giving a formal look. The environment is a room with wooden walls, and a large map or blueprint hangs on the wall in the background. The lighting highlights the three individuals, [1] [2] [4], and the focus is on them standing in a formal setting. [1] and [2] appear to be closer together, engaged in the setting’s activity, with [4] seeming to join or rejoin the group. [3] is no longer visible in the current frame. [4] has appeared, standing to the left side of [1] and [2]. [1] and [2] remain in similar positions as in the previous frame, but the group now includes [4].
[Uncaptioned image] timestamp$<$8$>$ 1: Person in a gray suit with a white shirt underneath. 2: Person wearing a white dress with long dark hair. 3: Person with long hair wearing a patterned dress, standing in the background. The current frame shows a group of three individuals indoors, with [1] on the left in a gray suit and white shirt, facing slightly towards [2], who is dressed in a white dress with long dark hair. [2] is looking at [1], suggesting an interaction or communication between them. [3] is slightly behind [2] and smiling, indicating a positive mood. The environment appears to be an office or meeting room with a large map or artwork on the wall in the background and a wooden wall, suggesting a formal or semi-formal setting. The lighting is bright, coming from the windows in the background, creating a clear but slightly shadowed detail on the individuals. From the previous frame to the current one, [1] and [2] appear to have shifted slightly closer to each other, with [2]’s head turned towards [1] indicating interaction. [3] is now visible in the scene, having entered from the right, which suggests a new addition to the group. [4] from the previous frame is no longer visible, indicating they may have exited the frame or moved out of view. The overall composition suggests a change in group dynamics as [3] enters and [1] and [2] interact more closely.
Table 9: Inst-IT Dataset Frame-level Annotation, Part II (frame 6-8). Please zoom in to view the instance ID labels.
Frame Instance-level Captions Image-level Captions Temporal Differences
[Uncaptioned image] timestamp$<$9$>$ 1: Wearing a light gray suit with a white shirt, standing with arms relaxed at the sides. 2: Wearing a sleeveless white dress, with black hair visible, standing sideways. 3: Clapping hands, wearing a dark, sleeveless floral-patterned dress. 4: Visible hands clapping, appearing on the left side of the frame. In the current frame, [1] is standing next to [2], both are positioned near a wooden wall, with a large framed picture or window in the background. [2] is wearing a white dress and stands slightly leaning towards [1], who is dressed in a gray suit. [3] is to the right, wearing a patterned dress and clapping her hands. On the left side of the frame, [4]’s hands are visible, indicating a clapping gesture. The environment appears to be well-lit, possibly indicating a celebratory or formal gathering. [4] has appeared in the current frame, clapping, which was not present in the previous frame. [1] and [2] have slightly shifted positions, indicating a minor adjustment in posture. The lighting in the room appears brighter in the current frame.
[Uncaptioned image] timestamp$<$10$>$ 1: [1] is wearing a grey suit with a white shirt. The person’s expression is neutral. 2: [2] is wearing a white dress, has long dark hair, and is smiling. 3: [3] is wearing a dark patterned dress, has long dark hair, and is smiling. 4: [4] is partially visible, clapping hands, wearing a long sleeve. In the current frame, [1] stands on the left wearing a grey suit and appears slightly more composed than before. [2], next to [1], in a white dress, continues smiling, directed towards [1]. [3] stands behind [2] with a continuous smile and hands still positioned as if clapping, indicating a joyous or celebratory mood. [4] is partially visible on the edge, with both hands shown as if engaged in clapping. The background remains the same, with wall decor and a wooden frame, suggesting an indoor setting. The lighting is consistent, highlighting a positive atmosphere. Between the previous and current frames, [1] has shifted from smiling to a neutral expression. [2]’s expression remains unchanged, still smiling. [3] continues to smile, maintaining the same engagement level. [4] shows hands in clapping motion slightly more forward than before. The physical positions of all individuals are largely the same, with slight adjustments in posture, possibly due to motion between shots.
[Uncaptioned image] timestamp$<$11$>$ 1: Individual in a grey suit with a light-colored shirt underneath. 2: Individual in a white dress with a flower in their hair. 3: Individual in a dark floral dress with bare shoulders. 4: Visible hand, partially in the frame, with a watch on the wrist. The current frame captures four adults in what appears to be an intimate celebration setting, inside a room with a wooden backdrop and a framed picture on the wall. [1] and [2] are the main focus, engaged in a kiss. Both are facing each other, with [1] in a grey suit and [2] in a white dress. [3] stands to the side, clapping, and appears joyous, indicating approval or celebration. The environment is that of a seemingly formal setting with elements suggesting a personal or official celebration. [4] is partially visible, with just a hand showing, suggesting a congratulatory gesture. Between the previous and current frames, [1] and [2] have moved from standing side by side to facing each other and kissing, indicating a change from a neutral to an intimate interaction. [3] continues to display a supportive gesture by clapping, suggesting this action started in the previous frame and continued into the current one. The position of [4] indicates movement from a neutral position to a congratulatory gesture, seen by the positioning of the arm and hand. The overall increase in physical interaction between [1] and [2] and the supportive gestures by [3] and [4] contribute to a more emotionally engaging scene in the current frame.
[Uncaptioned image] timestamp$<$12$>$ 1: Adult wearing a light grey suit with a white shirt. Short dark hair, clean-shaven, and standing upright. 2: Adult in a white, sleeveless dress. Long dark hair pulled back. Appears to be smiling with eyes partially closed. 3: Adult in a dark floral dress with a sleeveless design. Long dark hair down and clapping. In the current frame, [1] and [2] stand close together in the center of the image. [1] is wearing a grey suit with a white shirt and appears to be speaking or smiling. [2], dressed in a white dress, is leaning slightly towards [1] with a content expression. [3] is on the right, wearing a dark floral dress and clapping, seemingly celebrating with [1] and [2]. The environment is indoors with a wooden wall and a large framed picture in the background. The overall mood is celebratory, suggesting an event or occasion has taken place. Compared to the previous frame, [1] and [2] were previously kissing, but now they are standing apart, with [2] leaning slightly towards [1]. [1] has shifted from facing [2] to facing slightly outward and appears to be speaking or smiling. [3] remains in the same position but continues clapping, indicating ongoing celebration. The celebratory mood persists, reflecting a continuation of the event captured in the previous frame.
Table 10: Inst-IT Dataset Frame-level Annotation, Part III (frame 9-12). Please zoom in to view the instance ID labels.
Question Answer
What change occurs with [1]’s expression between $<$10$>$ and the previous frame? [1] changes from smiling to a neutral expression.
What activity are [1] and [2] involved in at $<$11$>$? [1] and [2] are engaged in a kiss.
What is the overall mood during $<$11$>$ as suggested by [3]’s actions? A celebratory or joyous event.
What interaction occurs between [1] and [2] at $<$5$>$? [1] holds [2]’s hand, suggesting an intimate gesture or exchange, likely a ring.
Who joins [1] and [2] in the frame at $<$7$>$? [4] appears in the frame, joining [1] and [2].
What changes in the group’s composition between $<$7$>$ and $<$8$>$? [3] reappears, and [4] is no longer visible.
What common setting element is seen throughout the frames $<$1$>$ to $<$12$>$? The scene is in an indoor setting with wooden paneling and framed art.
What type of event is likely taking place based on the atmosphere in $<$4$>$ and $<$6$>$? A formal event, possibly a wedding or official gathering.
What new elements are introduced in the scene at $<$2$>$? [5] holds a cellphone in the background, partially visible.
What is the mood and lighting like at $<$6$>$? The mood is formal and celebratory, with bright lighting enhancing this atmosphere.
What new background element appears at $<$7$>$? There is a map or blueprint on the wall.
What is notable about [5]’s actions at $<$3$>$? [5] is lifting an object above their head, possibly a piece of paper.
What is the setting like in $<$3$>$? The group is gathered near a wooden conference table in a formal setting.
How are [1] and [2] interacting at $<$8$>$? They are engaged in conversation or communication, indicated by body language and focus.
What does [1]’s expression suggest at $<$12$>$? [1] speaks or smiles, suggesting engagement with [2] or others.
What shift occurs in the focus of the camera between $<$5$>$ and $<$6$>$? The camera focuses more on individuals standing together, reducing focus on the foreground objects.
What are [3] and [4] doing at $<$9$>$? They clapping their hands in celebration.
What decorative element is visible at $<$2$>$? A bouquet of flowers lies on the table near [2].
How has the posture of [1] and [2] changed by $<$6$>$? [1] and [2] face slightly outward, suggesting a pose for a photograph or audience.
What overall physical change occurs between [1] and [2] from $<$10$>$ to $<$11$>$? There’s a noticeable increase in their physical interaction, enhancing emotional engagement.
Table 11: Inst-IT Dataset Open-Ended Question-Answer Pairs. The images of each frame are illustrated in Tab. 8 to Tab. 10.